What is the force needed to accelerate a 0.170 kg hockey puck from rest to 65.2 m/s over a period of 0.610 seconds?



1
Expert's answer
2020-12-23T10:33:23-0500

Given:

"m=0.170\:\rm kg"

"v_i=0\:\rm m/s, \; v_f=65.2\:\rm m/s"

"t=0.610\:\rm s"


The required force

"F=ma=m\times\frac{v_f-v_i}{t}\\
=0.170\times\frac{65.2-0}{0.610}=18.2\:\rm N"